OpenStudy (anonymous):

Simplify: ( x^2-16) /(2x^2-7x-15) divided by (x+4) /(4x^2-22x+10)

OpenStudy (anonymous):

try first factorizing each of the polynomials then change the division into a multiplication by flipping the second fraction cancel common terms

OpenStudy (anonymous):

Can you show me

OpenStudy (anonymous):

x^2-16=x^2-4^2=(x-4)(x+4) (difference of 2 squares) 2x^2-7x-15=2x^2-10x+3x-15=2x(x-5)+3(x-5)=(2x+3)(x-5) (grouping method) 4x^2-22x+10=4x^2-20x-2x+10=4x(x-5)-2(x-5)=(4x-2)x-5) now just rearrange and cancel common terms

OpenStudy (anonymous):

Can show me in fraction form

OpenStudy (anonymous):

good